OpenStudy (anonymous):

Can someone point me in the right direction with integrating 21-(e^0.03x) dx?

OpenStudy (anonymous):

you know the integration of 21, which is 21x.. the integration of e^(0.03x)dx= (1/0.03) e^(0.03x) =100/3 e^(0.03x) the final result is \[21x-100/3 e ^{0.03x}+c\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome